# A deterministic multiple-shift lattice algorithm for function approximation in Korobov and half-period Cosine spaces

### 01 Apr 2026 v1

@@ Abstract
Approximating multivariate periodic functions in weighted Korobov spaces via rank-1 lattices is bottlenecked by frequency aliasing. Existing optimal-rate methods rely on randomized constructions or massive pre-computations. We propose a fully deterministic multiple-shift lattice algorithm without pre pre-computations. First, we develop a uniform shift framework for aliased frequency fibers that decouples and reduces sampling costs. Second, leveraging the Chinese Remainder Theorem and the Weil bound, we introduce an adaptive hybrid construction that algebraically guarantees the full rank and bounded condition number of the reconstruction matrix. We rigorously prove that this deterministic method maintains the optimal convergence rate in the worst-case setting.
Furthermore, we extend this framework to non-periodic domains via the tent transformation. By establishing a strict projection equivalence, we prove the algorithm attains optimal $L_{2}$ and $L_{\infty}$ approximation orders in the half-period cosine space, successfully resolving an open theoretical problem posed by Suryanarayana et al. (2016). This mathematically validates the proposed algorithm as a generic meshless spectral solver for high-dimensional boundary value problems, such as the Poisson equation with Neumann conditions. Numerical experiments corroborate the theoretical bounds, demonstrating an order-of-magnitude reduction in sampling complexity over probabilistic baselines while ensuring absolute deterministic stability.
Keywords: Function approximation, Korobov space, Cosine space, deterministic construction, lattice rule.
@@ 2.3 Construction of Generating Vector
Corollary 1.
If $M^{1/\alpha}>4$ and $d=1$, then $R=1$.
Proof.
By Lemma 1, if $M^{1/\alpha}>4$, then $2^{m-1}>2$, i.e., $m>2$. Consider $d=1$, we have
$R_{\alpha,\bm{\gamma},N}(\bm{g})\leq 2\frac{(1+m/2)^{d}}{m}=\frac{2+m}{m}=1+\frac{2}{m}<2.$
Hence, $R_{\alpha,\bm{\gamma},N}(\bm{g})=1$. ∎
@@
The geometric implication of $R=1$ is that each aliasing fiber contains exactly one frequency from the hyperbolic cross $A_{\alpha,\bm{\gamma},M}$. In this trivial scenario, no frequency aliasing occurs within the target set, rendering a standard single rank-1 lattice rule (without any shifts) sufficient for exact reconstruction. Since practical numerical approximation tasks typically demand a large sample size $N$ such that the condition $M^{1/\alpha}>4$ is trivially satisfied, the one-dimensional case ($d=1$) naturally reduces to this non-aliasing setting. Consequently we focus on $d\geq 2$ and $R\geq 2$.
@@ 4.1 Existence of Good Shifts: A Probabilistic Baseline
To ensure a unique solution of (6), we require that the Gram matrix $G_{j}=B_{j}^{H}B_{j}$ for each fiber is invertible. As established in Corollary 1 and Remark 1, the trivial case where the fiber length $v=1$ implies no aliasing. In such a scenario, the Gram matrix reduces to the scalar $S$, and the system is trivially non-singular by assigning a single unshifted lattice ($S=1$ and $\bm{y}_{1}=\bm{0}$). Given our global assumption $d\geq 2$, we henceforth focus exclusively on the non-trivial case where $v\geq 2$.
@@
For a fixed $j=1,2,\ldots,J$, let $G_{j}=\sum_{s=1}^{S}X_{s}$ where $X_{s}=\bm{b}_{s}\bm{b}_{s}^{H}$ and $\bm{b}_{s}=[e^{-2\pi\mathrm{i}\bm{\ell}_{1}\cdot\bm{y}_{s}},\dots,e^{-2\pi\mathrm{i}\bm{\ell}_{v}\cdot\bm{y}_{s}}]^{\top}.$ Since $\bm{y}_{s}$ is uniformly distributed, $\mathbb{E}[X_{s}]=I_{v}$. Furthermore, $\lambda_{\max}(X_{s})=\|\bm{b}_{s}\|_{2}^{2}=v$ and $\lambda_{\min}(X_{s})=0$. Defining the zero-mean matrix $Z_{s}=I_{v}-X_{s}$, we have $\lambda_{\max}(Z_{s})\leq 1$ and $\mathbb{E}[Z_{s}^{2}]=(v-1)I_{v}$. Thus, the variance parameter is bounded by $\sigma^{2}=\|\sum_{s=1}^{S}\mathbb{E}[Z_{s}^{2}]\|_{2}=S(v-1)\leq S(R-1)$.
@@
$\mathbb{P}(\lambda_{\min}(G_{j}-SI_{v})\leq-\epsilon)=\mathbb{P}\left(\lambda_{\max}\left(\sum_{s=1}^{S}Z_{s}\right)\geq\epsilon\right)\leq v\cdot\exp\left(\frac{-\epsilon^{2}/2}{S(R-1)+\epsilon/3}\right)$
@@
  • Proposed Uniform Strategy: By sharing the shift across the fiber, the required total number of shifts is $S_{\text{new}}=\lceil 2KR\log N/t^{2}\rceil$ with an improved success probability of $1-N^{1-K}$.
Our uniform shift formulation achieves a rigorous reduction in sampling complexity from quadratic $\mathcal{O}(R^{2}\log N)$ to linear $\mathcal{O}(R\log N)$ with respect to the maximum fiber length $R$, while simultaneously removing the $R^{2}$ degradation from the probability bound.
@@ 5 The deterministic construction of multiple shifts
We now provide an explicit construction of multiple shifts $\{\bm{y}_{s}\}_{s=1}^{S}$ that guarantees the invertibility of the Gram matrix $G_{j}$ for each $j=1,\dots,J$. The entries of $G_{j}\in\mathbb{C}^{v_{j}\times v_{j}}$ are given by $G_{i,i}=S$ and, for $i\neq i^{\prime}$,
@@
Let $\bm{h}=\bm{\ell}_{i^{\prime}}-\bm{\ell}_{i}\neq\bm{0}$. By Gershgorin’s circle theorem, every eigenvalue $\lambda$ of $G_{j}$ lies in the disk:
@@
To ensure the non-singularity of $G_{j}$, it suffices to construct $\{\bm{y}_{s}\}_{s=1}^{S}$ such that the exponential sum in (8) is strictly bounded.
@@ 5.2 Deterministic Construction via the Union of Rank-1 Lattices with a Common Generator
$\bm{h}_{i}\cdot\bm{z}=\sum_{l=1}^{j-1}h_{i,l}z_{l}+h_{i,j}z_{j}+\underbrace{\sum_{l=j+1}^{d}h_{i,l}z_{l}}_{=0\text{ since }\bm{h}_{i}\in H_{j}}\neq 0\implies z_{j}\neq-\frac{1}{h_{i,j}}\sum_{l=1}^{j-1}h_{i,l}z_{l}.$
@@ 5.3 Adaptive Deterministic Strategy
  • Guaranteed Numerical Stability: The condition number analysis (Figure 2, right) explicitly verifies the dynamic transition mechanics. In low dimensions ($d<10$), $S_{\mathrm{adap}}$ selects the Polynomial curve, yielding a safely bounded condition number ($\kappa\approx 2.6$) strictly within the theoretical limit $\frac{1+t}{1-t}=39$. As $d\geq 10$, the strategy systematically switches to the Single-Lattice construction. This physically renders the columns of the system matrix perfectly orthogonal ($B^{H}B=pI_{v}$), driving the condition number down to the absolute optimal minimum of $\kappa(G_{j})=1$. The deterministic Adaptive framework thus achieves an order-of-magnitude reduction in sampling cost while mathematically cementing absolute numerical stability.
@@ 8.1 Periodic Test Function
Figure 4: Approximation errors for the periodic test function $f_{1}$ across dimensions $d=2$ and $d=4$. The left column demonstrates the spatial resolution accuracy bounded by $N$, while the right column illustrates the practical computational efficiency bounded by $N_{tot}=N\times S$. The reference lines indicate the optimal theoretical asymptotic slopes.

### 03 Apr 2026 v3

@@ Abstract
Approximating multivariate periodic functions in weighted Korobov spaces via rank-1 lattices is fundamentally limited by frequency aliasing. Existing optimal-rate methods rely on randomized constructions or large pre-computations. We propose a fully deterministic multiple-shift lattice algorithm without pre-computation. First, we develop a simplified multiple shift framework for aliased frequency fibers that reduces sampling costs. Second, leveraging the Chinese Remainder Theorem and the Weil bound, we introduce an adaptive hybrid construction that algebraically guarantees the full rank and bounded condition number of the reconstruction matrix. We rigorously prove that this deterministic method maintains the optimal convergence rate in the worst-case setting.
Furthermore, we extend this framework to non-periodic, half-period cosine spaces via the tent transformation. By establishing a strict projection equivalence, we prove that the algorithm attains optimal $L_{2}$ and $L_{\infty}$ approximation orders in the half-period cosine space, successfully resolving an open theoretical problem posed by Suryanarayana et al. (2016). This mathematically also validates the proposed algorithm as a generic meshless spectral solver for high-dimensional boundary value problems, such as the Poisson equation with Neumann conditions. Numerical experiments corroborate the theoretical bounds, demonstrating an order-of-magnitude reduction in sampling complexity over probabilistic baselines while ensuring absolute deterministic stability.
@@ 2.3 Construction of Generating Vector
For $d=1$, the CBC construction trivially yields the generator $g_{1}=1$. By our structural design, the truncation parameter $M$ is chosen such that the frequency set cardinality satisfies $|A_{\alpha,\bm{\gamma},M}|\leq N$, which ensures that for any distinct $k_{1},k_{2}\in A_{\alpha,\bm{\gamma},M}$, their difference satisfies $|k_{1}-k_{2}|<N$. Consequently, the aliasing condition $k_{1}-k_{2}\equiv 0\pmod{N}$ is impossible, yielding a maximum fiber length of $R=1$. Consequently, we focus on $d\geq 2$ and $R\geq 2$ throughout the remainder of this paper.
@@ 4.1 Existence of Good Shifts: A Probabilistic Baseline
To ensure a unique solution of (6), we require that the Gram matrix $G_{j}=B_{j}^{H}B_{j}$ for each fiber is invertible. As established in Remark 1, the trivial case where the maximum fiber length is $R=1$ implies an absolute absence of frequency aliasing. In such a scenario, the Gram matrix collapses to the scalar $S$, and the system is trivially non-singular by assigning a single unshifted lattice ($S=1$ and $\bm{y}_{1}=\bm{0}$). Given our global assumption $d\geq 2$, we henceforth focus exclusively on the non-trivial case where $R\geq 2$.
@@
For a fixed $j=1,2,\ldots,J$, let $G_{j}=\sum_{s=1}^{S}X_{s}$ where $X_{s}=\bm{b}_{s}\bm{b}_{s}^{H}$ and $\bm{b}_{s}=[e^{-2\pi\mathrm{i}\bm{\ell}_{1}\cdot\bm{y}_{s}},\dots,e^{-2\pi\mathrm{i}\bm{\ell}_{v}\cdot\bm{y}_{s}}]^{\top}.$ Since $\bm{y}_{s}$ is uniformly distributed, $\mathbb{E}[X_{s}]=I_{v_{j}}$. Furthermore, $\lambda_{\max}(X_{s})=\|\bm{b}_{s}\|_{2}^{2}={v_{j}}$ and $\lambda_{\min}(X_{s})=0$. Defining the zero-mean matrix $Z_{s}=I_{v_{j}}-X_{s}$, we have $\lambda_{\max}(Z_{s})\leq 1$ and $\mathbb{E}[Z_{s}^{2}]=({v_{j}}-1)I_{v_{j}}$. Thus, the variance parameter is bounded by $\sigma^{2}=\|\sum_{s=1}^{S}\mathbb{E}[Z_{s}^{2}]\|_{2}=S({v_{j}}-1)\leq S(R-1)$.
@@
$\mathbb{P}(\lambda_{\min}(G_{j}-SI_{v_{j}})\leq-\epsilon)=\mathbb{P}\left(\lambda_{\max}\left(\sum_{s=1}^{S}Z_{s}\right)\geq\epsilon\right)\leq{v_{j}}\cdot\exp\left(\frac{-\epsilon^{2}/2}{S(R-1)+\epsilon/3}\right)$
@@
Proposed Uniform Strategy: By sharing the shift across the fiber, the required total number of shifts is $S_{\text{new}}=\lceil 2KR\log N/t^{2}\rceil$ with an improved success probability of $1-N^{1-K}$. Our uniform shift formulation achieves a rigorous reduction in sampling complexity from quadratic $\mathcal{O}(R^{2}\log N)$ to linear $\mathcal{O}(R\log N)$ with respect to the maximum fiber length $R$, while simultaneously removing the $R^{2}$ degradation from the probability bound.
@@ 5 The deterministic construction of multiple shifts
We now provide an explicit construction of multiple shifts $\{\bm{y}_{s}\}_{s=1}^{S}$ that guarantees the invertibility of the Gram matrix $G_{j}$ for each $j=1,\dots,J$. The entries of $G=G_{j}\in\mathbb{C}^{v_{j}\times v_{j}}$ are given by $G_{i,i}=S$ and, for $i\neq i^{\prime}$,
@@
Let $\bm{h}=\bm{\ell}_{i^{\prime}}-\bm{\ell}_{i}\neq\bm{0}$. By Gershgorin’s circle theorem, every eigenvalue $\lambda$ of $G$ lies in the disk:
@@
To ensure the non-singularity of $G$, it suffices to construct $\{\bm{y}_{s}\}_{s=1}^{S}$ such that the exponential sum in (8) is strictly bounded.
@@ 5.2 Deterministic Construction via the Union of Rank-1 Lattices with a Common Generator
$\bm{h}_{i}\cdot\bm{z}=\sum_{l=1}^{j-1}h_{i,l}z_{l}+h_{i,j}z_{j}+\sum_{l=j+1}^{d}h_{i,l}z_{l}\neq 0\implies z_{j}\neq-\frac{1}{h_{i,j}}\sum_{l=1}^{j-1}h_{i,l}z_{l}.$
@@ 5.3 Adaptive Deterministic Strategy
Guaranteed Numerical Stability: The condition number analysis (Figure 2, right) explicitly verifies the dynamic transition mechanics. In low dimensions ($d<10$), $S_{\mathrm{adap}}$ selects the Polynomial curve, yielding a safely bounded condition number ($\kappa\approx 2.6$) strictly within the theoretical limit $\frac{1+t}{1-t}=39$. As $d\geq 10$, the strategy systematically switches to the Single-Lattice construction. This physically renders the columns of the system matrix perfectly orthogonal ($B^{H}B=pI_{v_{j}}$), driving the condition number down to the absolute optimal minimum of $\kappa(G_{j})=1$. The deterministic Adaptive framework thus achieves an order-of-magnitude reduction in sampling cost while mathematically cementing absolute numerical stability.
@@ 8.1 Periodic Test Function
Figure 4: The $L_{2}$ and $L_{\infty}$ errors for the periodic test function $f_{1}$ across dimensions $d=2$ and $d=4$.